# Safe Navigation in Uncertain Crowded Environments Using Risk Adaptive CVaR Barrier Functions

### v3

@@ I Introduction
Recently, [5, 6, 4, 7] have explored combining Conditional Value-at-Risk (CVaR) with Control Barrier Functions (CBFs) to achieve probabilistic safety, thereby providing a trade-off between conservatism and efficiency. In their frameworks, CBFs [8] act as safety filters that regulate control inputs, while CVaR [9] quantifies the expected risk under a given risk level. However, existing methods based on the CVaR barrier function (CVaR-BF) primarily rely on a fixed risk level that remains constant throughout the trajectory. This can limit their applicability in dynamic environments with crowded obstacles. We highlight that such a fixed risk level is not flexible enough: A low risk tolerance enhances safety but can render the optimization infeasible, whereas a high risk tolerance improves feasibility at the expense of safety.
To address this limitation, we propose a risk-adaptive navigation approach, a novel extension of the CVaR-BF framework that dynamically adjusts the risk level to maintain safety while ensuring trajectory feasibility. Instead of relying on a pre-specified risk parameter, our adaptive controller initializes the system with a conservative risk level (e.g., zero) and incrementally increases it until an optimization-feasible solution is achieved. In highly dynamic scenarios, where obstacles move unpredictably and rapidly, the robot requires sufficient time and space to respond and adjust its risk level. However, overly conservative strategies can limit feasible solutions in crowded environments [10, 11, 12], so an approach that maintains safety without excessively restricting the decision space is essential. To this end, we introduce the concept of a “Dynamic Zone”, where the original safety distance is adaptively expanded based on the relative position and velocity between the robot and its surrounding obstacles. The robot adjusts its trajectory before nearing obstacles, but only when needed to avoid unnecessary conservatism, while also extending the risk-adaptive range.
@@ II-2 Dynamic Obstacle Avoidance
For safety-critical navigation, a rigorous theoretical framework and analysis are required. In [17], a coupled planning approache is proposed to model the obstacles’ motion and solve a joint optimization problem with explicit safety constraints. However, coupled methods falter when the obstacle model is inaccurate [18], and accurately modeling obstacle motion is challenging. CBFs [8] have been developed as safety filters that mitigate unsafe control inputs. Some approaches partition the solution space into convex regions using separating hyperplanes [19, 4]. For example, [4] reformulates the collision avoidance problem by computing safe halfspaces based on dynamic obstacle sample trajectories. Recent work has proposed a CBF approach combined with velocity obstacles [10, 11, 12], where the velocity obstacle principle allows the vehicle to maintain forward motion without compromising safety. However, this method tends to trigger avoidance maneuvers even when obstacle are far away, thereby leading to an conservative behavior.
@@ V-A Dynamic Zone–Based Barrier Function
In contrast, functions based on velocity obstacles or collision cones [10, 11, 12] incorporate the relative motion between the robot and the obstacle. A typical candidate is:
@@ V-C Risk Adaptive CVaR-BF Optimization
Remark 2.
Feasibility condition: By setting $\beta=1$, Problem 1 reduces to the deterministic CBF-based optimization with multiple constraints and input limits. Therefore, as long as the deterministic problem is feasible, Problem 1 is always guaranteed to be feasible by setting $\beta=1$, even if no feasible solution exists for any $\beta<1$.

### v4

@@ I Introduction
Recently, [5, 4, 6] have explored combining Conditional Value-at-Risk (CVaR) with Control Barrier Functions (CBFs) to achieve probabilistic safety, thereby providing a trade-off between conservatism and efficiency. In their frameworks, CBFs [7] act as safety filters that regulate control inputs, while CVaR [8] quantifies the expected risk under a given risk level. However, existing methods based on the CVaR barrier function (CVaR-BF) primarily rely on a fixed risk level that remains constant throughout the trajectory. This can limit their applicability in dynamic environments with crowded obstacles. We highlight that such a fixed risk level is not flexible enough: A low risk tolerance enhances safety but can render the optimization infeasible, whereas a high risk tolerance improves feasibility at the expense of safety.
To address this limitation, we propose a risk-adaptive navigation approach, a novel extension of the CVaR-BF framework that dynamically adjusts the risk level to maintain safety while ensuring trajectory feasibility. Instead of relying on a pre-specified risk parameter, our adaptive controller initializes the system with a conservative risk level (e.g., zero) and incrementally increases it until an optimization-feasible solution is achieved. In highly dynamic scenarios, where obstacles move unpredictably and rapidly, the robot requires sufficient time and space to respond and adjust its risk level. However, overly conservative strategies can limit feasible solutions in crowded environments [9, 10], so an approach that maintains safety without excessively restricting the decision space is essential. To this end, we introduce the concept of a “Dynamic Zone”, where the original safety distance is adaptively expanded based on the relative position and velocity between the robot and its surrounding obstacles. The robot adjusts its trajectory before nearing obstacles, but only when needed to avoid unnecessary conservatism, while also extending the risk-adaptive range.
@@ II-2 Dynamic Obstacle Avoidance
For safety-critical navigation, a rigorous theoretical framework and analysis are required. In [15], a coupled planning approache is proposed to model the obstacles’ motion and solve a joint optimization problem with explicit safety constraints. However, coupled methods falter when the obstacle model is inaccurate [16], and accurately modeling obstacle motion is challenging. CBFs [7] have been developed as safety filters that mitigate unsafe control inputs. Some approaches partition the solution space into convex regions using separating hyperplanes [17, 4]. For example, [4] reformulates the collision avoidance problem by computing safe halfspaces based on dynamic obstacle sample trajectories. Recent work has proposed a CBF approach combined with velocity obstacles [9, 10], where the velocity obstacle principle allows the vehicle to maintain forward motion without compromising safety. However, this method tends to trigger avoidance maneuvers even when obstacle are far away, thereby leading to an conservative behavior.
@@ V-A Dynamic Zone–Based Barrier Function
In contrast, functions based on velocity obstacles or collision cones [9, 10] incorporate the relative motion between the robot and the obstacle. A typical candidate is: